The numbers 1998, 2997, 3996, 4995, ..., 8991 all have two distinct primes in common in their prime factorization. Find the sum of these two primes.
Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



Those numbers are all multiples of 999, which is 27*37 = 3*3*3*37.

The two distinct primes in the factorization of each of the numbers are 3 and 37.

ANSWER: 3+37 = 40


Answer by ikleyn(52797)

.

In order to facilitate your finding of the two distinct primes,  notice that the given numbers
represent an arithmetic progression with the common difference of

            2997 - 1998 = 999 = 3*333 = 3*3*111 = 3*3*3*37.

Next,  since the number  1998  is divisible by  3  and by  37  (check it directly !),
it implies that the given numbers  ALL  have  3  and  37  as their prime common divisors,
and  DO  NOT  HAVE  ANY  OTHER  prime common divisors.

Hence,  the problem wants you find the sum of  3  and  37,  which is   3+37 = 40.

ANSWER.   The sum of these two primes is  40.


Solved.